So far this yearThe average tactile during dinner service at a restaurant is $52 with a standard deviation of $11 to reach a sales goal, The next 25 checks during dinner service must average at least $50 each what is the probability this restaurant reaches it it’s sales goal? Write your answers as a decimal rounded to three places
1
Expert's answer
2020-07-23T18:17:19-0400

s="\sigma/\sqrt n=11/5=2.2"

P(x>50)=1-P(x<50)

P(x<50)

z="(x-\mu)/s" =-2/2.2=-0.91

using table, P=0.18141

P(x>50)=1-0.18141=0.819